For any integers n, let ® n be defined as the sum of the distinct prime factors of n. For instance, ® 36=5, because 2 and 3 are the only prime factors of 36 and
2+3=5.
What is the smallest value of w for which ® w=12 ?

Found 2 solutions by richard1234, solver91311:
Answer by richard1234(7193)

The only way to express 12 as a sum of distinct prime numbers is 5+7. Therefore the smallest value of n that satisfies is (5^1)(7^1) = 35, because 5+7 = 12.

Answer by solver91311(24713)



Hint: 5 + 7 = 12


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism